Item 5 How many excess electrons must be present on each sphere if the magnitude of the force of repulsion between them is 4.57102wiens? I Review Two small spheres spaced 20.0 centimeters apart have equal charge electrons Submit Reguest Answer

This is what I did, but my answer did not come out right.ざ 0 F5710N Af= 1.0.7 75 x 10-16 (1.6x10の -3 x l
Answer #1

let magnitude of charge be q

force between the charges is given by

F=9*109*q2/r2 = 9*109*q2/0.22=2.25*1011q2

Given force = 4.57*10-21 N

2.25*1011q2 = 4.57*10-21

q=1.42*10-16 C

charge on each electron = 1.6*10-19 C

number of excess electron on each sphere = 1.42*10-16/1.6*10-19 = 887